A solution contains 19 g of NaCl in a total volume of 325 mL. What is the NaCl concentration of the solution in g/L, % w/v and mol/L?


1
Expert's answer
2021-03-25T04:09:48-0400

Molar mass of NaCL = 58.44


Moles = "\frac{19}{58.44}=0.3251"


325 mL = 0.325L


Morality = "\frac{0.3251}{0.325}=1.0003 g/L"




1.0003×100= 100.03%w/v or 1.0003mol/L